# references.bib

### v1

@@ Abstract
Accurate estimates of Above Ground Biomass (AGB) are essential in addressing two of humanity’s biggest challenges, climate change and biodiversity loss. Existing datasets for AGB estimation from satellite imagery are limited. Either they focus on specific, local regions at high resolution, or they offer global coverage at low resolution. There is a need for a machine learning-ready, globally representative, high-resolution benchmark. Our findings indicate significant variability in biomass estimates across different vegetation types, emphasizing the necessity for a dataset that accurately captures global diversity. To address these gaps, we introduce a comprehensive new dataset that is globally distributed, covers a range of vegetation types, and spans several years. This dataset combines AGB reference data from the GEDI mission with data from Sentinel-2 and PALSAR-2 imagery. Additionally, it includes pre-processed high-level features such as a dense canopy height map, an elevation map, and a land-cover classification map. We also produce a dense, high-resolution (101010\,10m) map of AGB predictions for the entire area covered by the dataset. Rigorously tested, our dataset is accompanied by several benchmark models and is publicly available. It can be easily accessed using a single line of code, offering a solid basis for efforts towards global AGB estimation. The GitHub repository github.com/ghjuliasialelli/AGBD serves as a one-stop shop for all code and data.
@@ 1 Introduction
Measuring Above Ground Biomass (AGB) is a crucial capability to combat biodiversity loss and address the ongoing climate crisis. Accurate AGB estimates enable the quantification of carbon stocks, which play a pivotal role in carbon offsetting schemes. Furthermore, AGB is correlated with various biodiversity metrics, providing valuable insights into the structure of biodiversity hotspots [VanceChalcraft2010, Sonkoly2019, Li2018]. Traditional methods for measuring AGB are based on field work, either via destructive sampling, where sample trees are cut and weighed; or with non-destructive techniques that use tree dimensions (like tree height and trunk diameter at breast height) in tailored regression equations. Increasingly, LiDAR campaigns are replacing manual measurements with remote sensing, becoming the preferred method. However, despite their precision, the high costs and time requirements limit their global scalability. As a result, previous work has mainly focused on localized biomass estimation. Launched in 2019, the NASA GEDI mission uses an on-orbit laser altimeter aboard the International Space Station to scan the globe between $51.6^{\circ}$ north and south. Calibrated on coincident airborne LiDAR data and ground plot field inventories, the mission provides sparse AGB estimates across its observation domain. These estimates represent the largest existing reference AGB dataset, which can be combined with Machine Learning (ML) methods and Remote Sensing (RS) data to advance AGB estimation across the globe. We provide a ML-ready dataset consisting of coincident GEDI AGB estimates and various RS data products. Although freely accessible, obtaining remotely sensed data globally is tedious, and the sheer scale of it is prohibitive for ML research. The corresponding global dataset would require approximately 70TB of storage, $\approx 60$ times more than the complete ImageNet. For biomass estimation that data will need extensive preprocessing, and it would be too large to repeatedly train and test within reasonable time during method development. To reduce the scale,but to the greatest possible extent maintain the representativeness of the dataset, we derive a subset of regions whose vegetation distribution emulates the global one. The complex interactions between AGB and various vegetation types [Chen2023] motivate this choice.
@@
To highlight the accessibility of our dataset, we offer a fully preprocessed version that is compatible with all major machine learning frameworks, such as TensorFlow and PyTorch. It is hosted on HuggingFace [lhoest-etal-2021-datasets] and can be downloaded and used with just the following lines of code:
@@
  • We offer a machine learning-ready, easily accessible, globally representative dataset comprising coincident Aboveground Biomass (AGB) estimates together with various remote sensing data products.
@@ 2 Background
In recent years, remote sensing (RS) and machine learning (ML) have advanced various mapping tasks, ranging from human population [Metzger23] to canopy height [Lang2023]. Biomass estimation also has seen significant research, with [Zhang2019, Araza2022] providing a review of existing AGB datasets and maps, supplemented in Table 1. Currently, global-scale AGB datasets suffer from low spatial resolution (at best 100 meters), while high-resolution datasets (10-30 meters) are limited to restricted geographic areas. This dichotomy forces researchers to strike a compromise between generalization across space and spatial resolution. Notably, [Araza2022] highlight the challenge of achieving consistent accuracy across continental and global-scale AGB maps, due to the absence of a comprehensive global reference dataset. Furthermore, [Leonhardt2022] underscore the necessity for a dataset featuring globally distributed biomass reference records to improve regional studies. Similarly, [Biomassters] advocate for the release of more openly accessible, deep learning-ready data. [Tolan2024] highlight another aspect of this issue when they apply a correction based on GEDI to their recent 1-meter resolution canopy height maps for São Paulo (Brazil) and California (USA), aiming to compensate for the limited geographic diversity of their reference data. Taken together, these observations point at a critical gap in the literature: there is a need for a robust, globally diverse dataset to support high-resolution AGB mapping.
@@
[Biomassters] UNet Finland 10m Localized NFI S2, SAR
[Liu2023] RF, SVM, CNN Southwest China 10m Localized Destructive AGB measurements S1, S2, SAR
[Puliti2020] RF Northern Norway 10m Localized NFI ArcticDEM, S2
[SiyuLiu2023] UNet Europe 30m Localized NFI PlanetScope, LiDAR
[Leonhardt2022] CGAN Northwestern USA 30m Localized LiDAR-derived maps Landsat-8, ALOS-2 PALSAR-2
[Xu2017] Maximum Entropy estimator Democratic Republic of Congo 100m Localized, low-resolution LiDAR samples and allometry Landsat, ALOS-2 PALSAR-2, SRTM
@@ 3 Dataset
In this section, we discuss the rationale for creating the dataset, detail its components, and describe the process of its generation. For detailed information on data formats and downloading instructions, please refer to the supplementary material and the project website: github.com/ghjuliasialelli/AGBD.
@@ 3.1 Vegetation types analysis
To correctly represent the world’s vegetation types, we turn to the Copernicus Global Land Service Dynamic Land Cover map [LC2019], which delivers a global map of the land cover at 100100100\,100m spatial resolution. It consists of a discrete classification system with 21 classes (following UN-FAO’s Land Cover Classification System) that differentiates between open and closed forest types, with the following sub-types: evergreen needle-leaved (ENL), deciduous needle-leaved (DNL), evergreen broad-leaved (EBL), deciduous broad-leaved (DBL), mixed type, and unknown type. Additional land cover classes include shrubland, herbaceous vegetation, herbaceous wetland, moss and lichen, bare or sparse vegetation, cropland, built-up areas, snow and ice, and permanent water bodies. We refer the reader to [LC2019] for detailed descriptions of each class. We base our analysis on the $2019$ iteration of the product. Note that while the CGLS-LC100 map is global, we only consider vegetation types that are found within the GEDI coverage, as only there we have reference AGB data. As a consequence, vegetation types found only outside of the GEDI coverage are not accounted for. These are: open and closed DNL only found in Siberia; and moss and lichen, mostly found in Northern Canada, Siberia and Greenland.
Figure 2: Per-biome distribution of GEDI AGBD values and CCI residuals.
Left: global scale, Right: subset scale.
@@
The ESA CCI map is the most recent and adequate global AGB map available. It is provided by the European Space Agency (ESA), through its Climate Change Initiative (CCI) Biomass project. Global maps of AGB are provided for five epochs (2010, 2017, 2018, 2019 and 2020) at 100100100\,100m resolution. We observe that this map performs variably across different vegetation types, as shown in Figure 2. We utilise CCI AGB estimates and GEDI footprints from the year $2019$ to analyze residuals across various vegetation types within the GEDI coverage. Our analysis revealed significant variances, with underestimations in moss and lichen and overestimations in closed evergreen broadleaf forests, compelling us to adopt a design that accounts for differences between vegetation types. In the absence of a truly global dataset and facing the logistical barrier of large-scale remote sensing data downloads, we argue for a reasonably-sized dataset that mirrors the global vegetation distribution. To that end we select the following regions: California (USA), Cuba, Austria, Greece, Nepal, Shaanxi (China), French Guiana, Paraguay, Ghana, Tanzania and New Zealand. They are strategically chosen for their geographic and ecological diversity, as illustrated in Figures 1 and 3. Both the vegetation types and the ESA CCI residuals within our subset are in fairly good agreement with the global ones (Figure 2).
@@ 3.2 Modalities
In the following we describe the data sources, how they were processed, and how we assembled them into the final dataset. The process is illustrated in fig. 4. A more detailed explanation of each individual data field can be found on the project website.
GEDI, an altimeter mission (2018-2023) on the International Space Station, produced high-resolution laser observations of the Earth’s 3D structure. It collected billions of cloud-free observations from 51.6°N to 51.6°S. We downloaded all GEDI L4A (v2.1) footprints for the regions of interest from the start of the mission until Dec. 31st 2020. We only kept biomass values in the range $[0,500]$ Mg/ha, as suggested by previous literature [Carreiras2017]; we only retain the "power" beams, as previous work [MTClem, Lahssini2022-fp] has found that they provide more reliable estimates than the "coverage" beams; and as suggested by the GEDI team, we discard footprints where sensitivity $<0.95$, l4_quality_flag $\neq 1$, and selected_algorithm $=10$.
@@
PALSAR-2 (P2), on board the ALOS-2 satellite, is a Synthetic Aperture Radar (SAR) that can capture images at night and through clouds, making it a 24-hour, all-weather technology. The 252525\,25m PALSAR-2 mosaic is a seamless global image in WGS84 (EPSG:4326) projection, created by merging strips of SAR imagery, featuring both HH and HV polarization backscatter. It uses a different gridding system from ESA, so for each year we download the necessary ALOS tiles, then mosaic, crop and reproject them to align with Sentinel-2 tiles. The provided digital numbers ($DN$) are linear backscatter amplitudes. Pre-processing involves converting DN values to $\gamma_{0}$ values in decibels via $\gamma^{0}=10\cdot log_{10}(DN^{2})-83$.
Digital Elevation Model (DEM). The ALOS Global Digital Surface Model (AW3D30) offers 30-meter resolution elevation data. As for the PALSAR-2 mosaic we downloaded, mosaicked, and reprojected the DEM to align with Sentinel-2 tiles.
Land Cover (LC). We downloaded the Copernicus Global Land Service Dynamic Land Cover map at 100-meter resolution, cropped and reprojected it to match Sentinel-2 tiles.
Canopy Height (CH). [Lang2023] provides source code for canopy height mapping from Sentinel-2 imagery, using probabilistic deep learning. The model estimates CH for each input Sentinel-2 L2A image and merges redundant predictions from different dates to create a yearly CH map for each Sentinel-2 tile, along with a map of the associated standard deviations (STD). We follow this procedure to obtain yearly CH maps for 2019 and 2020 for all Sentinel-2 tiles within our regions of interest.
@@ 3.3 Dataset Assembly and Train/Test Split
To create an analysis-ready dataset, we crop all data into patches as follows: first all data sources are upsampled to a uniform resolution of 101010\,10m, using bilinear interpolation for continuous variables and nearest neighbor for categorical variables. We then iterate over all GEDI footprints within the area of interest, crop all raster data to 25$\times$25 pixel squares centered at the GEDI footprint, and store them in a HDF5 file. For more details please refer to the Appendix.
The splitting into training, validation and test sets is done not at the level of GEDI footprints but along Sentinel-2 tiles, meaning that all patches within the same $100\times 100\,$km2 tile will belong to the same set. To align with our geographical distribution and vegetation types, we split each country/region into training, validation and test portions and aim to match the distribution of vegetation types within each portion to that of the whole region. To that end we randomly explore the combinatorial assignment of tiles to the three portions until a sufficiently close match has been found. In each region $65\%$ of the tiles are used for training, $15\%$ for validation, and the remaining $20\%$ for testing. The detailed assignment per tile can be found on our project website.
4 Benchmark
@@ 4.1 Models
Early works on biomass estimation employed linear regression models, but the relationship between satellite obervations and AGB is more complex and can be better modeled with high-capacity ML methods [Li2020, Liu2023]. We provide a suite of pertinent ML models as benchmarks for the task.
@@
For the following deep learning methods, we always use a patch-wise training procedure: each patch (of size $12k+1\times 2k+$ 1) has one ground-truth pixel, its center. The model emits a prediction for each pixel in the patch, but only the central pixel prediction contributes to the loss. The optimisation of the network parameters is done in the standard way with mini-batch stochastic gradient descent, using the ADAM [ADAM] variant and batch size 256. Each model is trained five times with different random seeds to account for stochastic variability, an we always report the mean RMSE and standard deviation over the five runs. All models run on consumer hardware, specifically we used NVIDIA GeForce RTX 2080 Ti GPUs available through the ETH high-performance cluster. The training pipeline is illustrated in Figure 4.
@@
Lang et al. ($13$M parameters). The architecture developed in [Lang2023] for CH estimation, but trained with standard $\ell_{2}$-loss (rather than the original log-likelihood loss). It consists of a series of residual blocks with separable convolutions without any downsampling. For details, refer to [Lang2023].
Figure 4: Visualization of all data sources (except geographic coordinates) along with our training pipeline. Note the sparse GEDI labels, which are upgraded to dense AGB maps by the prediction.

### v2

@@ Abstract
Accurate estimates of Above Ground Biomass (AGB) are essential in addressing two of humanity’s biggest challenges: climate change and biodiversity loss. Existing datasets for AGB estimation from satellite imagery are limited. Either they focus on specific, local regions at high resolution, or they offer global coverage at low resolution. There is a need for a machine learning-ready, globally representative, high-resolution benchmark dataset. Our findings indicate significant variability in biomass estimates across different vegetation types, emphasizing the necessity for a dataset that accurately captures global diversity. To address these gaps, we introduce a comprehensive new dataset that is globally distributed, covers a range of vegetation types, and spans several years. This dataset combines AGB reference data from the GEDI mission with data from Sentinel-2 and PALSAR-2 imagery. Additionally, it includes pre-processed high-level features such as a dense canopy height map, an elevation map, and a land-cover classification map. We also produce a dense, high-resolution (101010\,10m) map of AGB predictions for the entire area covered by the dataset. Rigorously tested, our dataset is accompanied by several benchmark models and is publicly available. It can be easily accessed using a single line of code, offering a solid basis for efforts towards global AGB estimation.
keywords:
Remote sensing, Biomass Estimation, Machine Learning, Dataset.
@@ 1 Introduction
Measuring Above Ground Biomass (AGB) is crucial to combat biodiversity loss and address the ongoing climate crisis. Accurate AGB estimates enable the quantification of carbon stocks, which play a pivotal role in carbon offsetting schemes. Furthermore, AGB is correlated with various biodiversity metrics, providing valuable insights into the structure of biodiversity hotspots [VanceChalcraft2010, Sonkoly2019, Li2018]. Traditional methods for measuring AGB are based on field work, either via destructive sampling, where sample trees are cut and weighed; or with non-destructive techniques that use tree dimensions (like tree height and trunk diameter at breast height) in tailored regression equations. Increasingly, LiDAR campaigns are replacing manual measurements with Remote Sensing (RS), becoming the preferred method. However, despite their precision, the high costs and time requirements limit their global scalability. As a result, previous work has mainly focused on localized AGB estimation. Launched in 2019, the NASA GEDI mission uses an on-orbit laser altimeter aboard the International Space Station to scan the globe between $51.6^{\circ}$ north and south. Calibrated on coincident airborne LiDAR data and ground plot field inventories, the mission provides sparse AGB estimates across its observation domain. These estimates represent the largest existing reference AGB dataset, which can be combined with Machine Learning (ML) methods and RS data to advance AGB estimation across the globe. We provide a ML-ready dataset consisting of coincident GEDI AGB estimates and various RS data products. Although freely accessible, obtaining remotely sensed data globally is tedious, and the sheer scale of it is prohibitive for ML research. The corresponding global dataset would require approximately 70TB of storage, $\approx 60$ times more than the complete ImageNet. For AGB estimation that data will need extensive preprocessing, and it would be too large to repeatedly train and test within reasonable time during method development. To reduce the scale, but to the greatest possible extent maintain the representativeness of the dataset, we derive a subset of regions whose vegetation distribution emulates the global one. The complex interactions between AGB and various vegetation types [Chen2023] motivate this choice.
@@
To highlight the accessibility of our dataset, we offer a fully preprocessed version that is compatible with all major ML frameworks, such as TensorFlow and PyTorch. It is hosted on HuggingFace [lhoest-etal-2021-datasets] and can be downloaded and used with just the following lines of code:
@@
  • We offer a machine learning-ready, easily accessible, globally representative dataset comprising coincident AGB estimates together with various remote sensing data products.
@@ 2 Background
In recent years, remote sensing (RS) and machine learning (ML) have advanced various mapping tasks, ranging from human population [Metzger23] to canopy height [Lang2023]. Biomass estimation also has seen significant research, with [Zhang2019, Araza2022] providing a review of existing AGB datasets and maps, supplemented by Table 1. Currently, global-scale AGB datasets suffer from low spatial resolution (at best 100 meters), while high-resolution datasets (10-30 meters) are limited to restricted geographic areas. This dichotomy forces researchers to strike a compromise between generalization across space and spatial resolution. Notably, [Araza2022] highlight the challenge of achieving consistent accuracy across continental and global-scale AGB maps, due to the absence of a comprehensive global reference dataset. Furthermore, [Leonhardt2022] underscore the necessity for a dataset featuring globally distributed biomass reference records to improve regional studies. Similarly, [Biomassters] advocate for the release of more openly accessible, deep learning-ready data. [Tolan2024] highlight another aspect of this issue when they apply a correction based on GEDI to their recent 1-meter resolution canopy height maps for São Paulo (Brazil) and California (USA), aiming to compensate for the limited geographic diversity of their reference data. Taken together, these observations point at a critical gap in the literature: there is a need for a robust, globally diverse dataset to support high-resolution AGB mapping.
@@
[Biomassters] U-TAE, Swin U-Net TRansformer, U-Net++ Finland 10m Localized Airborne LiDAR and allometry S1, S2
[Liu2023] RF, SVM, CNN Southwest China 10m Localized Destructive AGB measurements S1, S2
[Puliti2020] RF Northern Norway 10m Localized Field measurements and allometry ArcticDEM, S2
[SiyuLiu2023] UNet Europe 30m Localized Field measurements, airborne LiDAR, allometry PlanetScope, LiDAR
[Leonhardt2022] CGAN Northwestern USA 30m Localized US Carbon Monitoring System AGB map Landsat-8, ALOS-2 PALSAR-2
[Xu2017] Maximum Entropy estimator Democratic Republic of Congo 100m Localized, low-resolution Airborne LiDAR and allometry Landsat-8, ALOS-2 PALSAR-2, SRTM
@@ 3 Dataset
In this section, we discuss the rationale for creating the dataset, detail its components, and describe the process of its generation. For detailed information on data formats and downloading instructions, please refer to the supplementary material and the to-be-released project website.
@@ 3.1 Vegetation types analysis
To correctly represent the world’s vegetation types, we turn to the Copernicus Global Land Service Dynamic Land Cover map [LC2019], which delivers a global map of the land cover at 100100100\,100m spatial resolution. It consists of a discrete classification system with 21 classes (following UN-FAO’s Land Cover Classification System) that differentiates between open (O) and closed (C) forest types, with the following sub-types: evergreen needle-leaved (ENL), deciduous needle-leaved (DNL), evergreen broad-leaved (EBL), deciduous broad-leaved (DBL), mixed type (MT), and unknown type (UT). Additional land cover classes include shrubland (Shrubs), herbaceous vegetation (HV), herbaceous wetland (HW), moss and lichen (Moss), bare or sparse vegetation (Bare), cropland (Crops), built-up areas (ignored), snow and ice (ignored), and permanent water bodies (ignored). We refer the reader to [LC2019] for detailed descriptions of each class. We base our analysis on the $2019$ iteration of the product (latest available). Note that while the CGLS-LC100 map is global, we only consider vegetation types that are found within the GEDI coverage, as only there we have reference AGB data. As a consequence, vegetation types found only outside of the GEDI coverage are not accounted for (the grey biomes in Figure 2). These are: open and closed DNL only found in Siberia; and moss and lichen, mostly found in Northern Canada, Siberia and Greenland.
Figure 2: Per-biome distribution of GEDI AGBD values and ESA CCI residuals. Gray biomes are outside of GEDI coverage.
@@
The ESA CCI map is the most recent and adequate global AGB map available. It is provided by the European Space Agency (ESA), through its Climate Change Initiative (CCI) Biomass project. Global maps of AGB are provided for five epochs (2010, 2017, 2018, 2019 and 2020) at 100100100\,100m resolution. We observe that this map performs variably across different vegetation types, as shown in Figure 2. We utilise CCI AGB estimates and GEDI footprints from the year $2019$ to analyze residuals across various vegetation types within the GEDI coverage. Our analysis revealed significant variances, with underestimations in moss and lichen and overestimations in closed evergreen broadleaf forests, compelling us to adopt a design that accounts for differences between vegetation types. In the absence of a truly global dataset and facing the logistical barrier of large-scale remote sensing data downloads, we argue for a reasonably-sized dataset that mirrors the global vegetation distribution. To that end we select the following regions: California (USA), Cuba, Austria, Greece, Nepal, Shaanxi (China), French Guiana, Paraguay, Ghana, Tanzania and New Zealand. They are strategically chosen for their geographic and ecological diversity, as illustrated in Figures 1 and 3. This selection was based on a careful process. We chose moderate-sized countries or large administrative regions to avoid areas that were too large or too small. We considered the number of available GEDI footprints, as regions further from the equator have higher revisit times, affecting data density. To ensure diverse ecological representation, we used the Shannon entropy and Simpson’s diversity index, based on land cover pixel values, to measure the biome diversity. Finally, regions were selected from various continents, with a balance above, below, and on the equator. Both the vegetation types and the ESA CCI residuals within our subset are in fairly good agreement with the global ones (Figure 2).
@@ 3.2 Modalities
In the following we describe the data sources, how they were processed, and how we assembled them into the final dataset. An overview is provided in Figure 4. A more detailed explanation of each individual data field can be found in the Appendix (Table 3 and Section 8.1).
GEDI, an altimeter mission (2018-2023) on the International Space Station, produced high-resolution laser observations of the Earth’s 3D structure. Each observation consists of a $25$m diameter footprint, with an associated AGB estimate. Due to its sampling pattern, GEDI has a (mostly) global yet sparse coverage: individual footprints from 51.6°N to 51.6°S. We downloaded all GEDI L4A[GEDI_L4A] (v2.1) footprints for the regions of interest from the start of the mission until Dec. 31st 2020. We only kept biomass values in the range $[0,500]$ Mg/ha, as suggested by previous literature [Carreiras2017]; we only retain the "power" beams, as previous work [MTClem, Lahssini2022-fp] has found that they provide more reliable estimates than the "coverage" beams; and as suggested by the GEDI team, we discard footprints where sensitivity $<0.95$, l4_quality_flag $\neq 1$, and selected_algorithm $=10$. Details on the process by which GEDI esimates the AGB can be found in the Appendix, in Section 8.4.
@@
PALSAR-2 (P2), on board the ALOS-2 satellite, is a Synthetic Aperture Radar (SAR) that can capture images at night and through clouds, making it a 24-hour, all-weather technology. The 252525\,25m PALSAR-2 yearly mosaic[Shimada_2014] is a seamless global image in WGS84 (EPSG:4326) projection, created by merging strips of SAR imagery for a year, featuring both HH and HV polarization backscatter. It uses a different gridding system from ESA, so for each year we download the necessary ALOS tiles, then mosaic, crop and reproject them to align with Sentinel-2 tiles. The provided digital numbers ($DN$) are linear backscatter amplitudes. Pre-processing involves converting DN values to $\gamma_{0}$ values in decibels via $\gamma^{0}=10\cdot log_{10}(DN^{2})-83$.
Digital Elevation Model (DEM). The ALOS Global Digital Surface Model (AW3D30)[Tadono_2014] offers 30-meter resolution elevation data. As for P2 mosaics, we downloaded, mosaicked, and reprojected the DEM to align with Sentinel-2 tiles.
Land Cover (LC). We downloaded the Copernicus Global Land Service Dynamic Land Cover map (Section 3.1) at 100-meter resolution, upsampled, cropped and reprojected it to match the Sentinel-2 tiles. The map includes two layers: the discrete classification, and the probability of the discrete classification (which acts as a quality indicator).
Canopy Height (CH). Lang et al. [Lang2023] provide source code for canopy height mapping from Sentinel-2 imagery, using probabilistic deep learning. The model estimates CH for each input Sentinel-2 L2A image and merges redundant predictions from different dates to create a yearly CH map for each Sentinel-2 tile, along with a map of the associated standard deviations (STD). We follow this procedure to obtain yearly CH maps for 2019 and 2020 for all Sentinel-2 tiles within our regions of interest.
Coordinates. We additionally provide latitude and longitude information to the model.
Figure 4: Visualization of all data sources along with our training pipeline. Note the sparse GEDI labels, which are upgraded to dense AGB maps by the prediction.
@@ 3.3 Dataset Assembly and Train/Test Split
To create an analysis-ready dataset, we crop all data into patches as follows. First, all data sources are up-sampled to a common resolution of 101010\,10m, using nearest neighbor for categorical variables (the S2 SCL band, and the LC probability band) and bilinear interpolation for continuous variables (all others). We then iterate over each GEDI footprint, and fetch the corresponding data: for S2, the closest product within the six months preceding the footprint acquisition; for the yearly available data products (CH and P2), the data from the year of the footprint acquisition; and the product itself for non time-dependent data products (DEM and LC). Once all data sources are temporally aligned, we crop all raster data to 25$\times$25 pixel squares centered on the GEDI footprint, and store them in a HDF5 file.
The splitting into training, validation and test sets is not done at the level of GEDI footprints, but at the level of the Sentinel-2 tiles, meaning that all patches within the same $100\times 100\,$km2 tile will belong to the same set. To align with our geographical distribution and vegetation types, we split each country/region into training, validation and test portions and aim to match the distribution of vegetation types within each portion to that of the whole region. To maintain our geographical distribution and vegetation types distribution, we perform a train/validation/test split for each region. We enforce that the vegetation types distribution within each region’s split should match the whole region’s distribution. This is done by randomly exploring the combinatorial assignment of tiles to the three splits until a sufficiently close match has been found. In each region $65\%$ of the tiles are used for training, $15\%$ for validation, and the remaining $20\%$ for testing. The detailed assignment per tile can be found on our project website.
4 Experiments
@@ 4.1 Models
Early works on biomass estimation employed linear regression models, but the relationship between satellite obervations and AGB is more complex and can be better modeled with high-capacity ML methods [Li2020, Liu2023]. We provide a selection of pertinent ML models as benchmarks for the task.
@@
For the following deep learning methods, we always use a patch-wise training procedure: each patch (of size $25\times 25$ or $15\times 15$) has one ground-truth pixel, its center. The model emits a prediction for each pixel in the patch, but only the central pixel prediction contributes to the loss. The optimisation of the network parameters is done in the standard way with mini-batch stochastic gradient descent, using the ADAM [ADAM] variant and batch size 256. Each model is trained five times with different random seeds to account for stochastic variability, and we report the mean RMSE and standard deviation over the five runs. All models run on consumer hardware, specifically we used NVIDIA GeForce RTX 2080 Ti GPUs available through our institution’s high-performance cluster. The training pipeline is illustrated in Figure 4.
@@
Lang et al. ($13$M parameters). The architecture developed in [Lang2023] for CH estimation, but trained (from scratch) for AGB estimation, with standard $\ell_{2}$-loss (rather than the original log-likelihood loss). It consists of a series of residual blocks with separable convolutions without any downsampling.